OpenStudy (anonymous):

Sandra makes a pizza.She puts mushrooms on 2 over 6 of the pizza.She adds green peppers to 5 over 8 of the pizza.Which topping covers more of the pizza?

OpenStudy (anonymous):

my sister is doing her math homework and she needs the steps through it.she`s in third grade

OpenStudy (anonymous):

So there are mushrooms on 1/3 of the pizza and peppers on 5/8 of the pizza. First find a common denominator.

OpenStudy (anonymous):

so 8 times 3 is 24. so thats the smallest common denominator you can have. so 1/3 is the same as 8/24, and 5/8 is the same as 15/24. so what is bigger, 8/24 or 15/24 ??

OpenStudy (anonymous):

You could always draw a picture as well

OpenStudy (anonymous):

good idea, but unfortunately you cannot draw on this website.

OpenStudy (anonymous):

dont you have to turn them to decimals to know

OpenStudy (anonymous):

he/she could, you could do fractions, decimals, a picture...lots of ways

OpenStudy (anonymous):

would you like to do decimals? fractions? or draw a picture? personally i think fractions is easier but it is your choice.

OpenStudy (anonymous):

could you teach me the fraction part

OpenStudy (anonymous):

Well an easy way to find common denominators is to multiply both denominators together: (3)(8) = 24 So 1/3 = (something)/24 and 5/8 = (something)/24 what times 3 will get you 24? Take that number and multiply it by both the numerator and the denominator: 1(8)/3(8) = 8/24 What times 8 will get you 24? 5(3)/8(3) = 15/24 So which fraction is bigger? 8/24 or 15/24 ?

OpenStudy (anonymous):

Well an easy way to find common denominators is to multiply both denominators together: (3)(8) = 24 So 1/3 = (something)/24 and 5/8 = (something)/24 what times 3 will get you 24? Take that number and multiply it by both the numerator and the denominator: 1(8)/3(8) = 8/24 What times 8 will get you 24? 5(3)/8(3) = 15/24 So which fraction is bigger? 8/24 or 15/24 ?

OpenStudy (anonymous):

Sorry about the double post...